OpenStudy (hashtag209):

A box contains a combination of solid-colored tickets: 1/10 of the tickets are green, 1/2 are red, 1/4 are blue, and the remaining 30 tickets are white. How many blue tickets are in the box? F. 10 G. 20 H. 50 J. 100 K. 200 I know the correct answer is H. 50 but I have no idea how they got that answer. I feel like the problem is missing key information in order to solve it, for example, how many tickets total were in the box. Can anyone explain to me how to get that answer, please?

OpenStudy (silverspeedfighter):

I think AceSpeedFighter Might Know this

OpenStudy (hashtag209):

I'm not really sure how to tag people. This is my first question, heh. How do you tag someone that might be able to help?

OpenStudy (silverspeedfighter):

hashtag i know you know who AceSpeedFighter is

OpenStudy (hashtag209):

Yes I do, but I thought for them to see the question you need to tag them. Am I wrong?

OpenStudy (silverspeedfighter):

@AceSpeedFighter

OpenStudy (hashtag209):

Thank you

OpenStudy (silverspeedfighter):

I would guess it woulb be D for my Di.....Dribble

OpenStudy (silverspeedfighter):

But thats just a guess

OpenStudy (silverspeedfighter):

Have you seen my name?

OpenStudy (hashtag209):

I'm not going to humor you, Silver. You're being condescending for no reason.

OpenStudy (silverspeedfighter):

SilverSpeedFighter-AceSpeedFighter-NovaSpeedFighter-SlySpeedFighter-SycSpeedFighter-SonicSpeedFighter-TrEcSpeedFighter

OpenStudy (hashtag209):

If you can't help me, then please don't reply just to be rude.

OpenStudy (silverspeedfighter):

How am i being rude? ;-;

OpenStudy (kropot72):

Let the total number of ticket be n. Now we can write: \[\large \frac{n}{10}+\frac{n}{2}+\frac{n}{4}=\frac{2n+10n+5n}{20}=\frac{17}{20}n\] Therefore the 30 white tickets must be 3/20 of the total number of tickets. \[\large \frac{3}{20}n=30\] So the total number of tickets is 200. The number of blue tickets is 200/4 = 50.

OpenStudy (silverspeedfighter):

I KNEW IT SEE

OpenStudy (silverspeedfighter):

Hashtag why do you think im being rude?

OpenStudy (silverspeedfighter):

i know you see this msg

OpenStudy (hashtag209):

Thank you Kropot. I appreciate the help.

OpenStudy (silverspeedfighter):

WOW

OpenStudy (silverspeedfighter):

Your being a real jerk

OpenStudy (silverspeedfighter):

Cmon stop being a jerk

OpenStudy (kropot72):

You're welcome :)